If the area of a triangle is 22 m squared and the base length is 5.5 m, what is the height?
Answer by nbedejeremiah(1)

Area = 22m^2
Base = 5.5m
A = 1/2 (b x h)
Using the formula above, 1/2 (5.5 x h) = 22
= 5.5h/2 = 22
Cross multiply (2 x 22 = 44)
5.5h = 44
Divide both sides by the coefficient of h.
5.5h/5.5 = 44/5.5
h = 8m